Use the market schedule below to answer the questions that follow:

P(¢) 5 9 13 17 21 25

       Qs 200 800 1400 Qd 2000 1700 1400

P = price, Qd = quantity demanded (in bags),

2000 2600 3200 1100 800 500

Qs = quantity supplied (in bags)

       a) Determine the equilibrium price and quantity.

b) If price is fixed at ¢21, calculate the excess demand/supply. c) If price falls from ¢13 to ¢9, calculate the:

i. price elasticity of demand. ii. arc price-elasticity of demand.


1
Expert's answer
2021-02-18T07:17:08-0500

(a) Equilibrium price and quantity are found where quantity supplied equals quantity demanded at the same price. From the table equilibrium price =13 and quantity = 1400bags.

(b) excess supply will be

"2600-1400=600"

Demand will reduce by

"800-1400=300"

(C) "e_{(d)}=\frac{dQ/Q}{dP/P}"

"=(1700-1400)\/1400"

"=0.21"

"21" %

"=(9-13)\/13=-0.31"

"31" %

"=21\/31=-0.697"

(I) "Arc Ed = [(Qd2 \u2013 Qd1) \/ midpoint Qd] \u00f7 [(P2 \u2013 P1) \/ midpoint P]"

Qd midpoint"=(1400+1700)\/2=1550"

P midpoint"=(13+9)\/2=11"

"=(1700-1400)1550=0.194"

"19.4%" %

"=(9-13)\/11=-0.364"

"36.4" %

"=19.4\/3.64=-0.53"